: Carbon dioxide CO2 is produced by burning fossil fuels such as oil and natural gas, and has been connected to global warming. The following table presents the average amounts (in metric tons) of

CO2 emissions for certain years per person in the United States and per person in the rest of the world.

Year Non-U.S. U.S.

1992

3.5

18.8

1993

3.3

19.7

1995

3.3

19.5

1996

3.4

19.7

2001

3.3

19.6

2002

3.3

19.6

2003

3.5

19.4

2004

3.7

19.6

2005

3.8

19.5

2006

3.7

19.0

Compute the least-squares regression line for predicting U.S. emissions from non-U.S. emissions. Round the slope and

y-intercept to at least four decimal places.
Answer #1

The following data are :

Non-U.S

U.S.

3.5

18.8

3.3

19.7

3.3

19.5

3.4

19.7

3.3

19.6

3.3

19.6

3.5

19.4

3.7

19.6

3.8

19.5

3.7

19.0

The independent variable is NonU.S, and the dependent variable is U.S.. In order to compute the regression coefficients, the following table needs to be used:

Non-U.S

U.S.

Non-U.S*U.S.

Non-U.S2

U.S.2

3.5

18.8

65.8

12.25

353.44

3.3

19.7

65.01

10.89

388.09

3.3

19.5

64.35

10.89

380.25

3.4

19.7

66.98

11.56

388.09

3.3

19.6

64.68

10.89

384.16

3.3

19.6

64.68

10.89

384.16

3.5

19.4

67.9

12.25

376.36

3.7

19.6

72.52

13.69

384.16

3.8

19.5

74.1

14.44

380.25

3.7

19.0

70.3

13.69

361

Sum =

34.8

194.4

676.32

121.44

3779.96

Based on the above table, the following is calculated:

\bar X = \frac{1}{n} \sum_{i=1}^{n} X_i = \frac{ 34.8}{ 10} = 3.48

\bar Y = \frac{1}{n} \sum_{i=1}^{n} Y_i = \frac{ 194.4}{ 10} = 19.44

SS_{XX} = \sum_{i=1}^{n} X_i^2 - \displaystyle\frac{1}{n}\left(\sum_{i=1}^{n} X_i\right)^2 = 121.44 - 34.8^2/10 = 0.33599999999997SS_{YY} = \sum_{i=1}^{n} Y_i^2 - \displaystyle\frac{1}{n}\left(\sum_{i=1}^{n} Y_i\right)^2 = 3779.96 - 194.4^2/10 = 0.82399999999961SS_{XY} = \sum_{i=1}^{n} X_i Y_i - \displaystyle\frac{1}{n}\left(\sum_{i=1}^{n} X_i\right) \left(\sum_{i=1}^{n} Y_i\right) = 676.32 - 34.8 \times 194.4/10 = -0.19200000000012

The regression coefficients (the slope m, and the y-intercept n) are obtained as follows:

m = \frac{SS_{XY}}{SS_{XX}} = \frac{ -0.19200000000012}{ 0.33599999999997} = -0.5714

n = \bar Y - \bar X \cdot m = 19.44 - 3.48 \times (-0.5714) = 21.4286

Therefore, we find that the regression equation is:

U.S. = 21.4286 - 0.5714 Non-U.S

please like)